OpenStudy (anonymous):

Simplify (x^9)^0 (x^7)^2

OpenStudy (solomonzelman):

\[(x^9)^0 (x^7)^2\]\[(x^9)^0=1\]\[ (x^7)^2~~~->~~~x^{14}\]

OpenStudy (anonymous):

any thing to the power of 0 is 1

OpenStudy (shamil98):

Anything to the power of zero is 1 , excluding zero. so (x^9)^ 0 = 1 1 * (x^7)^2 \[\huge (x^m)^n = x^ {m*n}\] so 1*x^14 = x^14

OpenStudy (solomonzelman):

@Jared01, do you want us to show you why is \[(ANYTHING)~^0=1\]

OpenStudy (solomonzelman):

\[s^0=s^{e-e}=s^e~/~s^e\]anything divided by itself is =1

OpenStudy (anonymous):

(\[\frac{ m ^{-1}m ^{5} }{ m ^{-2} }\]

OpenStudy (anonymous):

i need help with that

OpenStudy (shamil98):

\[\huge x^n * x^m = x^{n+m}\]

OpenStudy (shamil98):

\[\huge m^{-1} * m^5 = m^{-1+5}\]

OpenStudy (shamil98):

after that you remember when you divide exponents you subtract. \[\huge \frac{ x^m }{ x^n } = x^{m-n}\]